OpenStudy (anonymous):

f(x) = x-3/8; g(x) = 8x + 3, find g(f(x)).

hero (hero):

\(g(f(x)) = 8\left(\dfrac{x - 3}{8}\right) + 3 = x - 3 + 3 = x\)

OpenStudy (anonymous):

Why did you just multiply it by 8 instead of the entire equation 8x+3?

hero (hero):

You mean why did I put \(\dfrac{x - 3}{8}\) in place of \(x\) in \(8x + 3\) ?

OpenStudy (shamil98):

The 8 cancels it out .. 8(x-3)/8 you can cancel the two of them out.

OpenStudy (anonymous):

Ohh I see what you did! I thought it was supposed to be the opposite way. Thanks!

OpenStudy (anonymous):

It's fine. I still understand what you were trying to do. Thanks again! :)